paul invested $6000 in two accounts. one pay 5% simple interest and other pays 3%. if the interestearned on the 3% investment is $60 more than the interest earned on the 5% investment, how much is investedat each rate?
Answer by checkley77(12844)

.03x=.05(6,000-x)+60
.03x=300-.05x+60
.03x+.05x=360
.08x=360
x=360/.08
x=4,500 invested @ 3%.
6,000-4,500=1,500 invested @ 5%.
Proof:
.03*4,500=.05*1,500+60
135=75+60
135=135